OpenStudy (anonymous):

Solve the equation 3(4x - 3) - 19 = 8x - 4

OpenStudy (anonymous):

3(4x-3)-19=8x-4 12x-9-19=8x-4 12x-28=8x-4 12x-8x-28+4=0 4x-24=0 4x=24 x=24/4 x=6